Harvard Square Tour

12:00 PM - 3:00 PM Come and tour Cambridge’s world famous Harvard Square. From over 15 bookstores to exotic clothes, jewelry, and music, it has nearly everything. Meet on Student Center steps Cancelled in rain $1.20

MIT Museum Tour

1:00 PM - 2:00 PM Sample MIT’s rich tradition in science, engineering, architecture, and hacking at the MIT museum. Bring your temporary ID. Meet on Student Center steps Meet at Student Center info desk in rain free Athletic Facilities Tour

2:00 PM - 3:00 PM

Come tour MIT’s atheletic facilities and see all the wonderful things they have to offer. Bring $20 if you want to buy an athletic card. Meet on Student Center Steps Meet by the Student Center info desk in rain free

MIT Libraries Tour

3:00 PM - 4:00 PM

Learn about MIT’s library system and how best to use it. Meet in Lobby 7 Same if rain free MIT Offices Tour

4:00 PM - 5:00 PM

We’ll show you how to find all the offices you will deal with during your stay here: Financial Aid, Bursar, Registrar, Cashier’s, Housing, etc. Meet in Lobby 7 Same if rain free

Top of Boston Tour

6:15 PM - 8:45 PM

See the dramatic sunset view of Boston from the tallest building in New England, the Hancock Tower. A great orientation to Boston. Meet on Student Center steps Cancelled in heavy rain $4.45